OpenStudy (anonymous):

I need help ASAP, because I am extremely confused, but I need the answer right away.

OpenStudy (anonymous):

is it the same as the last one? i answered it there completely

OpenStudy (anonymous):

Evaluate sin(arccos(-9/sqrt(145)) a. 8/sqrt(145) b. -sqrt(145)/sqrt(248) c. sqrt(145)/sqrt(248) d. -8/sqrt(145)

OpenStudy (anonymous):

@satellite73 thank you!!!!!!!!!!!!!! :)

OpenStudy (anonymous):

third side is \(\sqrt{\sqrt{145}^2-9^2}=\sqrt{154-81}=\sqrt{64}=8\)

OpenStudy (anonymous):

answer is \(\frac{8}{\sqrt{145}}\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thank you so so so much!!!!!!!!!!!!!! :)